The rectangle has a perimeter of 40 centimeters.The length of the rectangle is 4 times as long as the width.Find the dimensions of the rectangle.
Answer by macston(5194)

.
W=width; L=length=4W; P=perimeter=40cm
P=2(L+W) Substitute for L
P=2(4W+W)
P=2(5W)
P=10W Divide each side by 10.
P/10=W Insert value for P
40cm/10=W
4cm=W ANSWER 1: The width is 4 cm.
L=4W=4(4cm)=16cm ANSWER 2: The length is 16 cm.
CHECK:
P=2(L+W)
40cm=2(16cm+4cm)
40cm=2(20cm)
40cm=40cm